# Extracting Sentence Embeddings from Pretrained Transformer Models

### v1

@@ 2.1 Composing word vectors
In the early work of [27], we can already find different weights for the members of the pairs of adjective-noun, noun-noun, or verb-object phrases. Later, work went from formal semantics to statistics-based measures, in particular, the tf-idf scores. They can be used as features on their own, such as in [58], but it is more wise to use them to weight neural word vectors. For example, the authors of [59] proposed the Composite Document Vector as a concatenation of idf-graded weighted word2vec vectors and tf-idf features. The authors of [60] proposed another better weighting approach, called Smooth Inverse Frequency (SIF). Here, higher-frequency words are down-weighted in a smooth fashion. The authors of [43] evaluated various compositional models and found that weighted averaging, in particular SIF, resulted in better performance in unsupervised similarity tasks that outperformed all other models. The author of [61] presented the uSIF method, an improvement to SIF, omitting hyperparameter tuning and constructing weighting with both word frequency and word vector length information. The authors of [62] also showed that the mean pooled output of transformers using tf-idf weights is better for clustering than the only regularly averaged output.
@@
A drastic case of weighting – a total removal of some words – is also shown to improve results. The authors of an older work [37] removed the stop words from the sequence before averaging. Meanwhile, as shown in [63, 15], the current performance of the BERT transformer model also is significantly improved in Semantic Text Similarity (STS) tasks if the most frequent tokens are removed. In other cases, the task itself focuses only on a few words, and others become redundant. The authors of [64] found that for the relation extraction task, some sentence embedding methods work better with shorter spans of words than the original sentences. They performed sentence segmentation in a way that a (sub)sentence would cover the identified entity mentions. Other authors of [64] also noted that SentenceBERT and Quickthought on spans or short segments containing two entity mentions are more clusterable than on the original sentences. These works also highlight the unequal contribution that each word has to the final sentence representation vector.
@@
Some works use more complex word weighting schemes. For example, the authors of [46] proposed using a form of attention to weight each hidden state of LSTM. Other authors of [65] proposed the idea that each word brings a novel orthogonal basis to the sentence. Therefore, the length of a projection in this direction can be converted into a word’s weight for use in the averaging. They come up with the final weight consisting of three scores: novelty, significance, and corpus-wise uniqueness. Authors of [64] found that for the methods analyzed, such weighting of GloVe embeddings made them the most clusterable. Similar work by [66] incorporated alignment and novelty scores and applied them to contextualized representations of SBERT. Here the alignment measures how well the word aligns with the neighboring ones, i.e., a well-aligned one is less informative and should be weighted less. The novelty, similar to [65], is expressed as the magnitude of the orthogonal component of the word to the subspace of neighboring words. Furthermore, the authors of [66] perform average weighting through all SBERT layers and also weight each layer-aggregated vector via $l_{1}$-normalized variance. This means that words that evolve faster across layers will get higher weights since they have greater variance. Although complex to implement, such methods are reported to provide some minor performance improvements.
@@
Surprisingly, the norm of word vectors has a large dispersion, as observed by the authors of [67, 68], and during the average pooling operation it acts as a weight of the word vector. The authors of [48] observed that their trained word vector in this way down-weights frequent tokens by itself, and this weighting follows the hypothesis of Luhn [69], a well-known information retrieval paradigm, stating that mid-rank terms are the most significant for discriminating content. Using these insights, the authors of [70] used the norm of word vectors as a proxy for the importance of words.
Clustering
@@
As shown by [71], averaging word vectors leads to a loss of information. It is in particular significant for longer text sequences, such as documents. Therefore, the idea was developed to aggregate vectors in a more smooth and information-preserving manner. The authors of [72] clustered the word vectors into $K$ groups using K-means. Then, for a document, each cluster vector is obtained as a sum of its constituent vectors. The final representation is then the concatenation of the cluster vectors and the inverse cluster frequency (icf) values, which are calculated using the idf values of the words present in the document. Other authors of [73] proposed a similar method, but viewed it as one that reduces the dimensionality from words to concepts. Concepts (as clusters) are created by clustering word vectors generated from word2vec, and frequencies of these concept clusters are used to represent document vectors. Finally, similarly to tf-idf applied to bag-of-words, a weighting scheme, concept frequency-inverse document frequency (cf-idf) is applied to acquired bag-of-concepts. The authors of [74] proposed the concepts of word containers and document containers to explain how such procedures bring benefits. Similarly to other works, they perform clustering of words into distinct clusters. Vectors belonging to the same cluster are averaged, while resulting representations among different clusters are concatenated. Using standard contextualized pre-trained transformer models with frozen weights, the authors of [75] instead of a simple mean pooling proposed to train a categorical variational autoencoder and reported the performance improvement on some STS tasks. This way, similar to clustering, the lower intermediate compressed representation would carry the essential representation of the sequence. Overall, using clustering of independent word vectors allows the document representation to be expressed in a smaller space of concepts whose concepts are otherwise erased due to destructive interference during the global averaging.
@@
The success of SCDV was supplemented by numerous other contributions. The authors of [77] instead of the Gaussian Mixture Models (GMM) used K-SVD (an algorithm for designing overcomplete dictionaries for sparse representation) [78], for the topic modeling. Furthermore, they used a newer word vector algorithm Doc2VecC and a modern SIF weighting and removal of the top principal component of [60]. Other authors of [79] with their SCDV-MS method additionally performed word sense disambiguation to dissect polysemous words into distinct vectors. In this way, the quality of the clusters improved. To induce sparsity, the authors also applied hard thresholding in an earlier stage in word cluster assignments and trained word vectors with Doc2vecC. Next, with the emergence of transformer models, the authors of [80] utilized contextualized representations in their SCDV+BERT(ctxd) method. First, the corpus is contextualized following the technique of [81]. That is, each corpus word is clustered among its individual recurrence contexts into distinct clusters corresponding to the different meanings of the same word. After such a procedure is repeated for all words, the vector for each word will now be the centroid vector of the closest cluster. The authors of [80] then repeated the original SCVD procedure and the result was an average improvement for STS tasks. Therefore, the idea of turning global averaging to local inside-cluster averaging did not change, only newer representation techniques and models were employed.
Spectral methods
@@
Text sequence comprising word embeddings can be thought of as a multidimensional signal over time. Therefore, temporal summarization techniques can be employed.
@@
One such method is the Discrete Cosine Transform (DCT), originally presented in [82]. It is an invertible function that maps an input sequence of $N$ real numbers to the coefficients of the $N$ number of orthogonal cosine basis functions. The DCT components are arranged in order of significance with the first one being proportional to the simple average of the sequence. DCT is used in data compression by preserving only the most important coefficients.
@@
In works of [83, 84] DCT was shown to outperform simple averaging of word embeddings. First, it can be attributed to the fact that DCT is structure-sensitive, as it captures signal dynamics. Second, the DCT has multiple coefficients, with the first one already corresponding to averaging, thus it should capture more information and enrich the performance of probing classifiers. The authors of [83] apply DCT along the sequence of words for each embedding dimension. Then, they retain lower-order coefficients and concatenate them to obtain overall feature patterns in the word sequence. One drawback of this technique is that short sentences must be padded with zeroed vectors. The following work of [84] also showed good results for DCT in multilingual and cross-lingual settings.
@@
In relation to the spectral decomposition of DCT, the authors of [85] proposed EigenSent that utilizes Higher Order Dynamic Mode Decomposition (HODMD). The method summarizes transitions in a sequence of words into one representative sequence embedding. The authors found that the best performance is achieved when such an embedding is concatenated with simple word vector averaging. This way, information on both the dynamics and the scale of the sequence is captured.
Using special tokens
@@
Modern transformer models have several special tokens that can play an important role during representation aggregation. For example, BERT has [CLS] designed for the next sentence prediction task, and it is supposed to carry the meaning of the input text sequence. The authors of [86] tried to use this property and trained the model to learn to aggregate everything in this [CLS] token. Other special tokens are [SEP], used to separate two text inputs, [MASK], to concentrate the representation of the masked word, padding, and various sentinel tokens (as in T5). Despite its intended purpose, special tokens are generally dismissed, as in [87, 14, 88], and a simple average is used instead. However, such tokens are important, as the authors of [89] after fine-tuning BERT found a clear tendency for earlier layers to pay attention to [CLS] and for later layers to [SEP].
@@
Recently, a prompt-based learning paradigm has emerged (see the survey [90]). Here, instead of fine-tuning the model to the downstream task, one reformulates it to look like the one that was being solved during the model pre-training process. Therefore, the main effort goes to prompt engineering, i.e., finding the most appropriate prompt for the task at hand. Such a strategy is in particular suitable for very large, even colossal language models that are difficult to train.
@@
The authors of [15] successfully used prompt engineering to better capture sentence representations from the BERT model. During the manual prompt search, the best prompt found was This sentence: "[X]" means [MASK] ., where [X] is replaced by the input sentence, and the output vector of token [MASK] is used as a final representation of the input sequence. Then, they further optimized the template by fine-tuning it on Natural Language Inference (NLI) data with the contrastive objective, while the BERT model weights were frozen. The authors used the continuous template technique of [91] where each template token is treated as a vector and optimized by gradient descent. The final template was shown to outperform multiple untrained baselines. The authors also showed that different templates can be used effectively to represent the same sentence with different points of view during supervised contrastive learning. The following work by [92] adapted the idea of using prompt instead of mean pooling and also used this representation extraction during contrastive learning.
Aggregating through layers
@@
One can imagine that “each layer will increasingly magnify small but significant differences” [44]. Therefore, the early layers capture more fundamental and low-level information [89], which dominates the learning of shallow lexical and meaning-related knowledge [93]. The authors of [75] reported that the word embeddings of the lower layers of BERT perform better than their upper layers on a word analogy task, and other authors of [53] showed that the first quarter of the models’ layers perform best in lexical semantic understanding. In the middle layers of the transformers, the hidden states are the most transferable [94] and contain the most relevant information [95]. The authors of [96, 97] report that the middle layers of the multilingual transformers are more multilingually-aligned. According to the authors of [93], with the inclusion of context in the upper layers, the encoded concepts evolve into a linguistic hierarchy where the middle and upper layers have a better representation of the core linguistic and semantic concepts. Finally, as the author of [98] explain, upper-layer representations become more context-specific.
@@
The general tendency, as shown by [99], is that the STS performance increases until the middle layers before decreasing toward the final ones. The authors of [99] even report that the final layer of most transformer models produces the worst-performing representations. Most scholars agree that this is due to the overspecialization of the last layers to the pre-training task [100]. Therefore, the final layers of BERT are the most task-specific [101] (but not as much as LSTM, as analyzed in [94]). Even during the fine-tuning phase, the authors of [89] determined that the last two layers encoded the highest share of task-specific features that are attributed to the score gain. Therefore, the last layers are specific to the training objective and cause problems if it has a large gap to the downstream task objective.
@@
However, some work in [102, 103] found that it was beneficial to apply post-processing to the final layers. This enabled them to restore the apparent drop in representational quality in later layers and gave large performance improvements. It is thought that anisotropy, known as the narrow cone of embeddings in vector space, may be to blame, which in [98] was found to increase from the earlier to the later layers. However, successful post-processing reveals that the last layers also possess relevant knowledge, but it is somehow obscured in the raw outputs.
@@
The question is how to compose a universal representation from the relevant pieces in multiple layers. One way is to try various combinations. The authors of [14] analyzed all possible two-layer combinations for BERT. They found that combining the top and bottom layers is better than using only the top layers. They tried to expand it to more layers but remained on a combination of 3 layers (L1+L2+L12). From the older works, we can mention the AdaSent model [104]. It operates on a hierarchy of representations derived from a pyramid-like multilayer network structure. The gating network is trained to adaptively select the most appropriate representation in the hierarchy for the task at hand. A similar gating system for dynamically deciding which intermediate transformer layers to use was proposed by the authors of [105].
@@
The contribution of each layer can also be accounted for by weighted average. The authors of ELMO [106], a contextualized model based on LSTM, present this configuration with additional weight parameters for each layer. Other authors investigating representations [107, 94, 95] call this weighting a scalar mix technique. The authors of [66] instead of learning layer weights for each word, compute them using complex alignment and novelty measures. For a few-layer model like ELMO, other authors of [108] found it useful to concatenate representations across layers while averaging across tokens, rather than using only the top layer. Other authors of [86] extract relevant information from intermediate hidden representations of BERT by treating them as positive samples during contrastive learning and determining which final sentence embedded in the tuned model should be close.
@@
In conclusion, there is no generally accepted method for taking advantage of representations across all layers in an unsupervised setting. In the other, supervised case, the whole network is fine-tuned, and the whole hierarchy of representations is taught to surface according to the task at hand.
Other means of composition

### v2

@@ 2.1.4 Weighted Average
In an early work [27], we can already find different weights for the members of the pairs of adjective–noun, noun–noun, or verb–object phrases. Later, work went from formal semantics to statistics-based measures, in particular the tf-idf scores. They can be used as features on their own, such as in [58], but it is better to use them to weight neural word vectors. For example, the authors of [59] proposed the Composite Document Vector as a concatenation of idf-graded weighted Word2Vec vectors and tf-idf features. The authors of [60] proposed another better weighting approach, called Smooth Inverse Frequency (SIF). Here, higher-frequency words are down-weighted smoothly. The authors of [43] evaluated various compositional models and found that weighted averaging, in particular SIF, resulted in better performance in unsupervised similarity tasks that outperformed all other models. The author of [61] presented the uSIF method, an improvement to SIF, omitting hyperparameter tuning and constructing weighting with both word frequency and word vector length information. The authors of [62] also showed that the mean pooled output of transformers using tf-idf weights is better for clustering than the only regularly averaged output.
@@
A drastic case of weighting—a total removal of some words—is also shown to improve results. The authors of an older work [37] removed the stop words from the sequence before averaging. Meanwhile, as shown in [63, 15], the current performance of the BERT transformer model is also significantly improved in Semantic Text Similarity (STS) tasks if the most frequent tokens are removed. In other cases, the task itself focuses only on a few words, and others become redundant. The authors of [64] found that for the relation extraction task, some sentence embedding methods work better with shorter spans of words than the original sentences. They performed sentence segmentation in a way that a (sub)sentence would cover the identified entity mentions. Other authors of [64] also noted that SentenceBERT and Quickthought on spans or short segments containing two entity mentions are more clusterable than on the original sentences. These works also highlight the unequal contribution that each word has to the final sentence representation vector.
@@
Some works use more complex word weighting schemes. For example, the authors of [46] proposed a form of attention to weight each hidden state of LSTM. Other authors of [65] proposed the idea that each word brings a novel orthogonal basis to the sentence. Therefore, the length of a projection in this direction can be converted into a word’s weight for use in the averaging. They come up with the final weight, consisting of three scores: novelty, significance, and corpus-wise uniqueness. Authors of [64] found that for the methods analyzed, such weighting of GloVe embeddings made them the most clusterable. Similar work by [66] incorporated alignment and novelty scores and applied them to contextualized representations of SBERT. Here, the alignment measures how well the word aligns with the neighboring ones, i.e., a well-aligned one is less informative and should be weighted less. The novelty, similar to [65], is expressed as the magnitude of the orthogonal component of the word to the subspace of neighboring words. Furthermore, the authors of [66] perform average weighting through all SBERT layers and also weight each layer-aggregated vector via $l_{1}$-normalized variance. This means that words that evolve faster across layers will receive higher weights since they have greater variance. Although complex to implement, such methods are reported to provide some minor performance improvements.
@@
Surprisingly, the norm of word vectors has a large dispersion, as observed by the authors of [67, 68], and during the average pooling operation, it acts as a weight of the word vector. The authors of [48] observed that their trained word vector in this way down-weights frequent tokens by itself, and this weighting follows the hypothesis of Luhn [69], a well-known information retrieval paradigm, stating that mid-rank terms are the most significant for discriminating content. Using these insights, the authors of [70] used the norm of word vectors as a proxy for the importance of words.
2.1.5 Clustering
@@
As shown by [71], averaging word vectors leads to a loss of information. It is in particular significant for longer text sequences, such as documents. Therefore, the idea was developed to aggregate vectors in a more smooth and information-preserving manner. The authors of [72] clustered the word vectors into $k$ groups using $k$-means. Then, for a document, each cluster vector is obtained as a sum of its constituent vectors. The final representation is then the concatenation of the cluster vectors and the inverse cluster frequency (icf) values, which are calculated using the idf values of the words present in the document. Other authors of [73] proposed a similar method, but viewed it as one that reduces the dimensionality from words to concepts. Concepts (as clusters) are created by clustering word vectors generated from Word2Vec, and frequencies of these concept clusters are used to represent document vectors. Finally, similarly to tf-idf applied to bag-of-words, a weighting scheme, concept frequency-inverse document frequency (cf-idf), is applied to acquired bag-of-concepts. The authors of [74] proposed the concepts of word containers and document containers to explain how such procedures bring benefits. Similarly to other works, they perform clustering of words into distinct clusters. Vectors belonging to the same cluster are averaged, while resulting representations among different clusters are concatenated. Using standard contextualized pre-trained transformer models with frozen weights, the authors of [75], instead of a simple mean pooling, proposed training a categorical variational autoencoder and reported the performance improvement on some STS tasks. This way, similar to clustering, the lower intermediate compressed representation would carry the essential representation of the sequence. Overall, using clustering of independent word vectors allows the document representation to be expressed in a smaller space of concepts, whose concepts are otherwise erased due to destructive interference during the global averaging.
@@
The success of SCDV was supplemented by numerous other contributions. The authors of [77], instead of the Gaussian Mixture Models (GMMs), used K-SVD (an algorithm for designing overcomplete dictionaries for sparse representation) [78] for the topic modeling. Furthermore, they used a newer word vector algorithm Doc2VecC and a modern SIF weighting and removal of the top principal component of [60]. Other authors of [79], with their SCDV-MS method, additionally performed word sense disambiguation to dissect polysemous words into distinct vectors. In this way, the quality of the clusters improved. To induce sparsity, the authors also applied hard thresholding in an earlier stage in word cluster assignments and trained word vectors with Doc2vecC. Next, with the emergence of transformer models, the authors of [80] utilized contextualized representations in their SCDV + BERT (ctxd) method. First, the corpus is contextualized following the technique of [81]. That is, each corpus word is clustered among its individual recurrence contexts into distinct clusters corresponding to the different meanings of the same word. After such a procedure is repeated for all words, the vector for each word will then be the centroid vector of the closest cluster. The authors of [80] then repeated the original SCVD procedure and the result was an average improvement for STS tasks. Therefore, the idea of turning global averaging to local inside-cluster averaging did not change; only newer representation techniques and models were employed.
2.1.6 Spectral Methods
@@
A text sequence comprising word embeddings can be interpreted as a multidimensional signal over time. Therefore, temporal summarization techniques can be employed.
@@
One such method is the Discrete Cosine Transform (DCT), originally presented in [82]. This invertible function maps an input sequence of $N$ real numbers to the coefficients of the $N$ number of orthogonal cosine basis functions. The DCT components are arranged in order of significance, with the first one being proportional to the simple average of the sequence. DCT is used in data compression by preserving only the most important coefficients.
@@
In [83, 84], DCT was shown to outperform simple averaging of word embeddings. First, it can be attributed to the fact that DCT is structure-sensitive, as it captures signal dynamics. Second, the DCT has multiple coefficients, with the first one already corresponding to averaging; thus, it should capture more information and enrich the performance of probing classifiers. The authors of [83] apply DCT along the sequence of words for each embedding dimension. Then, they retain lower-order coefficients and concatenate them to obtain overall feature patterns in the word sequence. One drawback of this technique is that short sentences must be padded with zeroed vectors. The following work [84] also showed good results for DCT in multilingual and cross-lingual settings.
@@
Concerning the spectral decomposition of DCT, the authors of [85] proposed EigenSent, which utilizes Higher-Order Dynamic Mode Decomposition (HODMD). The method summarizes transitions in a sequence of words into one representative sequence embedding. The authors found that the best performance is achieved when such an embedding is concatenated with simple word vector averaging. This way, information on both the dynamics and the scale of the sequence is captured.
2.1.7 Using Special Tokens
@@
Modern transformer models have several special tokens that can play an important role during representation aggregation. For example, BERT has [CLS] designed for the next sentence prediction task, and it is supposed to carry the meaning of the input text sequence. The authors of [86] tried to use this property and trained the model to learn to aggregate everything in this [CLS] token. Other special tokens are [SEP], used to separate two text inputs, [MASK], to concentrate the representation of the masked word, padding, and various sentinel tokens (as in T5). Despite its intended purpose, special tokens are generally dismissed, as in [87, 14, 88], and a simple average is used instead. However, such tokens are important, as the authors of [89], after fine-tuning BERT, found a clear tendency for earlier layers to pay attention to [CLS] and for later layers to pay attention to [SEP].
@@
Recently, a prompt-based learning paradigm has emerged (see the survey [90]). Here, instead of fine-tuning the model to the downstream task, one reformulates it to look like the one that was being solved during the model pre-training process. Therefore, the main effort goes to prompt engineering, i.e., finding the most appropriate prompt for the given task. Such a strategy is especially suitable for very large, even colossal language models that are difficult to train.
@@
The authors of [15] successfully used prompt engineering to better capture sentence embeddings with the BERT model. During the manual prompt search, the best prompt found was “This sentence: "[X]" means [MASK].”, where [X] is replaced by the input sentence, and the output vector of token [MASK] is used as a final representation of the input sequence. Then, they further optimized the template by fine-tuning it on Natural Language Inference (NLI) data with the contrastive objective, while the BERT model weights were frozen. The authors used the continuous template technique of [91], where each template token is treated as a vector and optimized by gradient descent. The final template was shown to outperform multiple untrained baselines. The authors also showed that different templates can be used effectively to represent the same sentence with different points of view during supervised contrastive learning. The following work by [92] adapted the idea of using prompt instead of mean pooling and also used this representation extraction during contrastive learning.
2.1.8 Aggregating through Layers
@@
One can imagine that “each layer will increasingly magnify small but significant differences” [44]. Therefore, the early layers capture more fundamental and low-level information [89], which dominates the learning of shallow lexical and meaning-related knowledge [93]. The authors of [75] reported that the word embeddings of the lower layers of BERT perform better than their upper layers on a word analogy task, and other authors of [53] showed that the first quarter of the models’ layers perform best in lexical semantic understanding. In the middle layers of the transformers, the hidden states are the most transferable [94] and contain the most relevant information [95]. The authors of [96, 97] report that the middle layers of the multilingual transformers are more multilingually aligned. According to the authors of [93], with the inclusion of context in the upper layers, the encoded concepts evolve into a linguistic hierarchy where the middle and upper layers have a better representation of the core linguistic and semantic concepts. Finally, as the author of [98] explains, upper-layer representations become more context-specific.
@@
A general tendency, as shown by [99], is that the STS performance increases until the middle layers before decreasing toward the final ones. The authors of [99] even report that the final layer of most transformer models produces the worst-performing representations. Most scholars agree that this is due to the overspecialization of the last layers to the pre-training task [100]. Therefore, the final layers of BERT are the most task-specific [101] (but not as much as LSTM, as analyzed in [94]). Even during the fine-tuning phase, the authors of [89] determined that the last two layers encoded the highest share of task-specific features attributed to the score gain. Therefore, the last layers are specific to the training objective and cause problems if it differs a lot from the downstream task objective.
@@
However, some work in [102, 103] found that it was beneficial to apply post-processing to the final layers. This enabled them to restore the apparent drop in representational quality in later layers and gave large performance improvements. It is thought that anisotropy, known as the narrow cone of embeddings in vector space, may be to blame, which in [98] was found to increase from the earlier to the later layers. However, successful post-processing reveals that the last layers also possess relevant knowledge, but it is somewhat obscured in the raw outputs.
@@
The question is how to compose a universal representation from the relevant pieces in multiple layers. One way is to try various combinations. The authors of [14] analyzed all possible two-layer combinations for BERT. They found that combining the top and bottom layers is better than using only the top layers. They tried to expand it to more layers but remained on a combination of three layers (L1 + L2 + L12). From the older works, we can mention the AdaSent model [104]. It operates on a hierarchy of representations derived from a pyramid-like multilayer network structure. The gating network is trained to adaptively select the most appropriate representation in the hierarchy for the given task. A similar gating system for dynamically deciding which intermediate transformer layers to use was proposed by the authors of [105].
@@
The contribution of each layer can also be accounted for by weighted average. The authors of ELMO [106], a contextualized model based on LSTM, present this configuration with additional weight parameters for each layer. Other authors investigating representations [107, 94, 95] call this weighting a scalar mix technique. The authors of [66], instead of learning layer weights for each word, compute them using complex alignment and novelty measures. For a few-layer model like ELMO, other authors of [108] found it useful to concatenate representations across layers while averaging across tokens, rather than using only the top layer. Other authors of [86] extract relevant information from intermediate hidden representations of BERT by treating them as positive samples during contrastive learning and determining which final sentence embedded in the tuned model should be close.
@@
In conclusion, there is no generally accepted method for taking advantage of representations across all layers in an unsupervised setting. In the other, supervised case, the whole network is fine-tuned, and the whole hierarchy of representations is taught to surface representations according to the task at hand.
2.1.9 Other Means of Composition